Suppose a subdivision on the southwest side of Denver, Colorado, contains 1,500 houses. The subdivision was built in 1983. A sample of 120 houses is selected randomly and evaluated by an appraiser. If the mean appraised value of a house in this subdivision for all houses is $229,000, with a standard deviation of $8,700, what is the probability that the sample average is greater than $230,500? Appendix A Statistical Tables. (Round the values of z to 2 decimal places. Round your answer to 4 decimal places.)
Answer #1

µ = 229000, σ = 8700, n = 120

P(X̅ > 230500) =

= P( (X̅-μ)/(σ/√n) > (230500-229000)/(8700/√120) )

= P(z > 1.89)

= 1 - P(z < 1.89)

Using excel function:

= 1 - NORM.S.DIST(1.89, 1)

= 0.0294